# Networked Communication for Mean-Field Games with Function Approximation and Empirical Mean-Field Estimation

### v1

@@ Abstract
Recent works have provided algorithms by which decentralised agents, which may be connected via a communication network, can learn equilibria in Mean-Field Games from a single, non-episodic run of the empirical system. However, these algorithms are given for tabular settings: this computationally limits the size of players’ observation space, meaning that the algorithms are not able to handle anything but small state spaces, nor to generalise beyond policies depending on the ego player’s state to so-called ‘population-dependent’ policies. We address this limitation by introducing function approximation to the existing setting, drawing on the Munchausen Online Mirror Descent method that has previously been employed only in finite-horizon, episodic, centralised settings. While this permits us to include the population’s mean-field distribution in the observation for each player’s policy, it is arguably unrealistic to assume that decentralised agents would have access to this global information: we therefore additionally provide new algorithms that allow agents to estimate the global empirical distribution based on a local neighbourhood, and to improve this estimate via communication over a given network. Our experiments showcase how the communication network allows decentralised agents to estimate the mean-field distribution for population-dependent policies, and that exchanging policy information helps networked agents to outperform both independent and even centralised agents in function-approximation settings, by an even greater margin than in tabular settings.
@@ 1 Introduction
The mean-field game (MFG) framework (Lasry and Lions 2007; Huang, Malhamé, and Caines 2006) can be used to address the difficulty faced by multi-agent reinforcement learning (MARL) regarding computational scalability as the number of agents increases. It models a representative agent as interacting not with other individual agents in the population on a per-agent basis, but instead with a distribution over the other agents, known as the mean field. The MFG framework analyses the limiting case when the population consists of an infinite number of symmetric and anonymous agents, that is, they have identical reward and transition functions which depend on the mean-field distribution rather than on the actions of specific other players. The solution to this game is the mean-field Nash equilibrium (MFNE), which can be used as an approximation for the Nash equilibrium (NE) in a finite-agent game, with the error in the solution reducing as the number of agents N tends to infinity (Saldi, Başar, and Raginsky 2018; Anahtarci, Kariksiz, and Saldi 2023; Yardim, Goldman, and He 2024; Toumi, Malhame, and Le Ny 2024; Hu and Zhang 2024). MFGs have thus been applied to a wide variety of real-world problems: see Laurière et al. (2022) for applications and references.
@@
Recent works argue that classical algorithms for solving MFGs rely on assumptions and methods that are likely to be undesirable in real-world applications, emphasising that desirable qualities for practical MFG algorithms include: learning from the empirical distribution of $N$ agents (without generation or manipulation of this distribution by the algorithm itself or an external oracle/simulator); learning from a single, continued system run that is not arbitrarily reset as in episodic learning; model-free learning; decentralisation; and fast practical convergence (Yardim et al. 2023; Benjamin and Abate 2024). While these works address these desiderata, they do so only in settings in which the state and action spaces are small enough that the Q-function can be represented by a table, limiting their approaches’ scalability.
@@
Moreover, in those works, as in many others on MFGs, agents only observe their local state as input to their Q-function (which defines their policy). This is sufficient when the solved MFG is expected to have a stationary distribution (‘stationary MFGs’) (Laurière et al. 2022; Xie et al. 2021; Anahtarci, Kariksiz, and Saldi 2023; uz Zaman et al. 2023; Yardim et al. 2023; Benjamin and Abate 2024). However, in reality there are numerous reasons for which agents may benefit from being able to respond to the current distribution. Recent work has thus increasingly focused on these more general settings where it is necessary for agents to have so-called ‘master policies’ (a.k.a. population-dependent policies) which depend on both the mean-field distribution and their local state (Cardaliaguet et al. 2015; Carmona, Delarue, and Lacker 2016; Perrin et al. 2022; Wu et al. 2024; Laurière et al. 2022; Lauriere et al. 2022; Laurière et al. 2022).
@@
The distribution is a high-dimensional (and thus large) observation object, and also takes a continuum of values. Therefore a population-dependent Q-function cannot be represented exactly in a table and must be approximated. To address these limitations while maintaining the desiderata for real-world applications given in recent works, we introduce function approximation to the MFG setting of decentralised agents learning from a single, non-episodic run of the empirical system, allowing this setting to handle larger state spaces and to accept the mean-field distribution as an observation input. To overcome the difficulties of training non-linear approximators in this context, we use the so-called ‘Munchausen’ trick, introduced by Vieillard, Pietquin, and Geist (2020) for single-agent RL, and extended to MFGs by Lauriere et al. (2022), and to MFGs with population-dependent policies by Wu et al. (2024).
@@
We particularly explore this in the context of networked communication between decentralised agents (Benjamin and Abate 2024). We demonstrate that this communication brings two specific benefits over the purely independent setting (while also removing the undesirable assumption of a centralised learner, which in the real world may be unrealistic, a computational bottleneck and a vulnerable single point of failure). Firstly, when the Q-function is approximated rather than exact, some updates lead to better performing policies than others. Allowing networked agents to propagate better performing policies through the population leads to faster practical convergence than in the purely independent-learning case, and often even than in the centralised case. Secondly, we argue that in the real world it is unrealistic to assume that decentralised agents, endowed with local state observations and limited (if any) communication radius, would be able to observe the global mean-field distribution and use it as input to their Q-functions / policy. We therefore further contribute two setting-dependent algorithms by which decentralised agents can estimate the global distribution from local observations, and further improve their estimates by communication with neighbours.
@@
We focus on ‘coordination games’, where agents can increase their individual rewards by following the same strategy as others and therefore have an incentive to communicate policies, even if the MFG setting itself is technically non-cooperative. Thus our work can be applied to real-world problems in e.g. traffic signal control, formation control in swarm robotics, and consensus and synchronisation e.g. for sensor networks. In summary, our contributions are:
@@
    • We use Munchausen RL for the first time in an infinite-horizon MFG context, unlike the finite-horizon settings in Lauriere et al. (2022) and Wu et al. (2024).
    • In doing so, we also present the first use of function approximation for solving MFGs from continued, non-episodic runs of the empirical system; this is unlike Yardim et al. (2023) and Benjamin and Abate (2024), which look at tabular settings.
@@
  • Rather than assuming that agents have access to this global knowledge, as in prior works, we present two novel algorithms that allow decentralised agents to locally estimate the empirical distribution and to improve these estimates by communication with others.
  • We present numerical results showcasing these algorithms, especially the two benefits of the decentralised communication scheme, which significantly outperforms both the independent and centralised settings.
The paper is structured as follows. Preliminaries are given in Sec. 2 and our core learning and policy improvement algorithm is presented in Sec. 3. Our mean-field estimation and communication algorithms are given in Sec. 4, and experiments are presented in Sec. 5. See Appx. B for a more detailed ‘Related work’ section. Ways to extend our algorithms are discussed in the ‘Future Work’ section (Appx. C).
@@ 2.1 Mean-field games
  • It is unrealistic to assume that decentralised agents with a possibly limited communication radius can observe the global mean-field distribution, so we allow agents to form a local estimate $\tilde{\hat{\mu}}^{i}_{t}$ which can be improved by communication with neighbours. Here we have $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$.
@@
Definition 1 (N-player symmetric anonymous games)
@@
Definition 2 (Induced mean-field flow)
@@
Definition 3 (Mean-field discounted return)
@@
Definition 4 (Best-response (BR) policy)
@@
Definition 5 (Mean-field Nash equilibrium (MFNE))
@@
  • $\pi^{*}$ is a best response to $\boldsymbol{\mu}^{*}$, i.e. $\pi^{*}\in BR(\boldsymbol{\mu})$;
@@
Previous works have shown that, in tabular settings, it is possible for a finite population of decentralised agents (each of which is permitted to have a distinct population-independent policy $\pi^{i}$) to learn the MFNE using only the empirical distribution $\hat{\mu}_{t}$, rather than the exactly calculated mean-field flow (Yardim et al. 2023; Benjamin and Abate 2024). This MFNE is in turn an approximate NE for the original finite-agent game. In this work we provide algorithms to perform this in non-tabular and population-dependent settings, and demonstrate them empirically.
@@
Instead of performing the computationally expensive process of finding a $BR$ at each iteration, we can use a form of policy iteration for MFGs called Online Mirror Descent (OMD). This involves beginning with an initial policy $\pi_{0}$, and then at each iteration $k$, evaluating the current policy $\pi_{k}$ with respect to its induced mean-field flow $\boldsymbol{\mu}=I(\pi_{k})$ to compute its Q-function $Q_{k+1}$. To stabilise the learning process, we then use a weighted sum over this and past Q-functions, and set $\pi_{k+1}$ to be the softmax over this weighted sum, i.e. $\pi_{k+1}(\cdot|s,\mu)=softmax\left(\frac{1}{\tau_{q}}\sum^{k}_{\kappa=0}{Q_{\kappa}}(s,\mu,\cdot)\right)$. $\tau_{q}$ is a temperature parameter that scales the entropy in Munchausen RL (Vieillard, Pietquin, and Geist 2020); note that this is a different temperature to the one agents use when selecting which communicated parameters to adopt, denoted $\tau^{comm}_{k}$ and discussed in Sec. 3.2.
If the Q-function is approximated non-linearly using neural networks, it is difficult to compute this weighted sum. The so-called ‘Munchausen trick’ addresses this by computing a single Q-function that mimics the weighted sum using implicit regularisation based on the Kullback-Leibler (KL) divergence between $\pi_{k}$ and $\pi_{k+1}$ (Vieillard, Pietquin, and Geist 2020). Using this reparametrisation gives Munchausen OMD (MOMD), detailed further in Sec. 3.1 (Lauriere et al. 2022; Wu et al. 2024). MOMD does not bias the MFNE, and has the same convergence guarantees as OMD (Hadikhanloo 2017; Pérolat et al. 2022a; Wu et al. 2024).
@@ 2.3 Networks
Definition 6 (Time-varying network)
@@
One of these networks $\mathcal{G}^{comm}_{t}$ defines which agents are able to communicate information to each other at time $t$. The second network $\mathcal{G}^{obs}_{t}$ is a graph defining which agents are able to observe each other’s states, which we use in general settings for estimating the mean-field distribution from local information. The structure of the two networks may be identical (e.g. if embodied agents can both observe the position (state) of, and exchange information with, other agents within a certain physical distance from themselves), or different (e.g. if agents can observe the positions of nearby agents, but only exchange information with agents by which they are connected via a telephone line, which may connect agents over long distances).
@@
Definition 7 (Time-varying state-visibility graph)
@@ 3.1 Q-network and update
0:  initial states {$s^{i}_{0}$}i, $i=1,\ldots,N$; $t\leftarrow 0$
1:  $\forall i$ : Randomly initialise parameters $\theta^{i}_{0}$ of Q-networks $\check{Q}_{\theta^{i}_{0}}(o^{i}_{t},\cdot)$, and set $\pi^{i}_{0}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{0}}(o,\cdot)\right)(a)$
@@
14:     $\forall i$ : Set $\pi^{i}_{k+1}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k+1}}(o,\cdot)\right)(a)$
15:     $\forall i:$ $\sigma^{i}_{k+1}=0$
@@
17:        Take step $\forall i:a^{i}_{t}\sim\pi^{i}_{k}(\cdot|o^{i}_{t}),r^{i}_{t}=R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t}),s^{i}_{t+1}\sim P(\cdot|s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})$; $t\leftarrow t+1$
18:        $\forall i:$ $\sigma^{i}_{k+1}=\sigma^{i}_{k+1}+\gamma^{e}(r^{i}_{t}+h(\pi^{i}_{k+1}(s^{i}_{t})))$
@@
23:        $\forall i:J^{i}_{t}\leftarrow\{j\in\mathcal{N}:(i,j)\in\mathcal{E}_{t}$}
@@
29:  return  policies $\pi^{i}_{K}$}i, $i=1,\ldots,N$
Lines 1-14 of our novel Alg. 1 contain the core Q-function/policy update method. Agent $i$ maintains a neural network parametrised by $\theta^{i}_{k}$ to approximate its Q-function: $\check{Q}_{\theta^{i}_{k}}(o,\cdot)$. The agent’s policy is given by $\pi_{\theta^{i}_{k}}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k}}(o,\cdot)\right)(a)$. When appropriate, we denote the policy $\pi^{i}_{k}(a|o)$ for simplicity. Each agent maintains a buffer (with size $M$) of collected transitions of the form $\left(o^{i}_{t},a_{t},r_{t},o^{i}_{t+1}\right)$. At each iteration $k$, they empty their buffer (Line 3) before collecting $M$ new transitions (Lines 4-7); each decentralised agent $i$ then trains its Q-network $\check{Q}_{\theta^{i}_{k}}$ via $L$ training updates as follows (Lines 8-12). For training purposes, $i$ also maintains a target network $\check{Q}_{\theta^{i,^{\prime}}_{k,l}}$ with the same architecture but parameters $\theta^{i,^{\prime}}_{k,l}$ copied from $\theta^{i}_{k,l}$ less regularly than $\theta^{i}_{k,l}$ themselves are updated, i.e. only every $\nu$ learning iterations (Line 11). At each iteration $l$, the agent samples a random batch $B^{i}_{k,l}$ of $|B|$ transitions from its buffer (Line 9). It then trains its neural network using stochastic gradient descent to minimise the following empirical loss (Line 10):
Definition 8 (Empirical loss for Q-network)
The empirical loss is given by

### v2

@@ Abstract
Recent algorithms allow decentralised agents, possibly connected via a communication network, to learn equilibria in Mean-Field Games from a non-episodic run of the empirical system. However, these algorithms are for tabular settings: this computationally limits the size of agents’ observation space, meaning the algorithms cannot handle anything but small state spaces, nor generalise beyond policies depending only on the agent’s local state to so-called ‘population-dependent’ policies. We address this limitation by introducing function approximation to the existing setting, drawing on the Munchausen Online Mirror Descent method that has previously been employed only in finite-horizon, episodic, centralised settings. While this permits us to include the mean field in the observation for players’ policies, it is unrealistic to assume decentralised agents have access to this global information: we therefore also provide new algorithms allowing agents to locally estimate the global empirical distribution, and to improve this estimate via inter-agent communication. We show theoretically that exchanging policy information helps networked agents outperform both independent and even centralised agents in function-approximation settings. Our experiments demonstrate this happening empirically, by an even greater margin than in tabular settings, and show that the communication network allows decentralised agents to estimate the mean field for population-dependent policies.
@@ 1 Introduction
The mean-field game (MFG) framework Lasry and Lions (2007); Huang et al. (2006) can be used to address the difficulty faced by multi-agent reinforcement learning (MARL) regarding computational scalability as the number of agents grows Yardim and He (2024); Zeng et al. (2024). It models a representative agent as interacting not with other individual agents in the population on a per-agent basis, but instead with a distribution over the other agents, called the mean field. The MFG framework analyses the limiting case when the population consists of an infinite number of symmetric and anonymous agents, that is, they have identical reward and transition functions which depend on the mean-field distribution rather than on the actions of specific other players. The solution to this game is the mean-field Nash equilibrium (MFNE), which can be used as an approximation for the Nash equilibrium (NE) in a finite-agent game (which is harder to solve in itself), with the error in the solution reducing as the number of agents N tends to infinity Saldi et al. (2018); Anahtarci et al. (2023); Yardim et al. (2024); Toumi et al. (2024); Hu and Zhang (2024); Chen et al. (2024). MFGs have thus been applied to a wide variety of real-world problems: see Laurière et al. (2022a) for examples.
@@
Recent works argue that classical algorithms for solving MFGs rely on assumptions and methods that are likely to be undesirable in real-world applications (e.g. swarm robotics or autonomous vehicles), emphasising that desirable qualities for practical MFG algorithms include: learning from the empirical distribution of $N$ agents (i.e. this distribution is generated only by the policies of the agents, rather than being updated by the algorithm itself or an external oracle/simulator); learning online from a single, non-episodic system run (i.e. similar to above, the population cannot be arbitrarily reset by an external controller); model-free learning; decentralisation; and fast practical convergence Yardim et al. (2023); Benjamin and Abate (2023). While these works address these desiderata, they do so only in settings in which the state and action spaces are small enough that the Q-function can be represented by a table, limiting their approaches’ scalability.
@@
Moreover, in those works, as in many others on MFGs, agents only observe their local state as input to their Q-function (which defines their policy). This is sufficient when the solved MFG is expected to have a stationary distribution (‘stationary MFGs’) Laurière et al. (2022a); Xie et al. (2021); Anahtarci et al. (2023); Zaman et al. (2023); Yardim et al. (2023); Benjamin and Abate (2023). However, in reality there are numerous reasons for which agents may benefit from being able to respond to the current distribution. Recent work has thus increasingly focused on these more general settings where it is necessary for agents to have so-called ‘master policies’ (a.k.a. population-dependent policies) which depend on both the mean-field distribution and their local state Cardaliaguet et al. (2015); Carmona et al. (2016); Perrin et al. (2022); Wu et al. (2024); Laurière et al. (2022a, b, a).
@@
The distribution is a large, high-dimensional observation object, taking a continuum of values. Therefore a population-dependent Q-function cannot be represented exactly in a table and must be approximated. To address these limitations while maintaining the desiderata for real-world applications given in recent works, we introduce function approximation to the MFG setting of decentralised agents learning online from a single, non-episodic run of the empirical system, allowing this setting to handle larger state spaces and to accept the mean-field distribution as an observation input.To overcome the difficulties of training non-linear approximators in this context, we use the so-called ‘Munchausen’ trick, introduced by Vieillard et al. (2020) for single-agent RL, and extended to MFGs by Laurière et al. (2022b), and to MFGs with population-dependent policies by Wu et al. (2024).
@@
We particularly explore this in the context of networked communication between decentralised agents Benjamin and Abate (2023). We demonstrate that communication brings two specific benefits over the purely independent setting (while also removing the undesirable assumption of a centralised learner, which in the real world may be unrealistic, a computational bottleneck and a vulnerable single point of failure). Firstly, when the Q-function is approximated rather than exact, some updates lead to better performing policies than others. Allowing networked agents to propagate better performing policies through the population (reminiscent of the use of fitness functions in distributed evolutionary algorithms Eiben and Smith (2015); Hart et al. (2015)) leads to faster learning than in the purely independent case and often even than in the centralised case, as we show both theoretically and empirically. Secondly, we argue that in the real world it is unrealistic to assume that decentralised agents, endowed with local state observations and limited (if any) communication radius, would be able to observe the global mean-field distribution and use it as input to their Q-functions / policy. We therefore further contribute two setting-dependent algorithms by which decentralised agents can estimate the global distribution from local observations, and further improve their estimates by communication with neighbours.
@@
We focus on ‘coordination’ games, where agents can increase their individual rewards by following the same strategy as others and therefore have an incentive to communicate policies, even if the MFG setting itself is technically non-cooperative. Thus our work can be applied to real-world problems in e.g. traffic signal control, formation control in swarm robotics, and consensus and synchronisation e.g. for sensor networks Soleimani et al. (2024). [^1] We further preempt concerns about the appropriateness of communication in competitive settings by wondering whether self-interested agents would be any more likely to want to obey a central learner as has usually been assumed. Moreover we show that self-interested communicating agents can obtain higher returns than independent agents even in non-coordination games (Fig. 5), indicating that they do have incentive to communicate. In summary, our contributions are:
@@
    • We use Munchausen RL for the first time in an infinite-horizon MFG context (for finite-horizon see Laurière et al. (2022b); Wu et al. (2024)).
    • This constitutes the first use of function approximation for solving MFGs from a single, non-episodic run of the empirical system (for tabular settings see Yardim et al. (2023); Benjamin and Abate (2023)).
@@
  • Rather than assuming that agents have access to this global knowledge as in prior works, we present two additional novel algorithms allowing decentralised agents to locally estimate the empirical distribution and to improve these estimates by inter-agent communication.
  • We show theoretically that networked agents may learn faster than both centralised and independent agents in the function-approximation setting.
  • We support this with extensive experiments, where our results showcase the two benefits of the decentralised communication scheme, which significantly outperforms both the independent and centralised settings.
The paper is structured as follows. We give preliminaries in Sec. 2 and our core learning and policy improvement algorithm in Sec. 3. We present our mean-field estimation and communication algorithms in Sec. 4, theoretical results in Sec. 5 and experiments in Sec. 6. Sup. Mat. D contains an extended ‘Related work’ section.
@@ 2.1 Mean-field games
  • It is unrealistic to assume that decentralised agents with a possibly limited communication radius can observe the global mean field, so we allow agents to form a local estimate $\tilde{\hat{\mu}}^{i}_{t}$ which can be improved by communication with neighbours. Here we have $o^{i}_{t}=(s^{i}_{t},\tilde{\hat{\mu}}^{i}_{t})$.
@@
Definition 1 (N-player symmetric anonymous games).
@@
Definition 2 (Induced mean-field flow).
@@
Definition 3 (Mean-field discounted return).
@@
Definition 4 (Best-response (BR) policy).
@@
Definition 5 (Mean-field Nash equilibrium (MFNE)).
@@
  • $\pi^{*}$ is a best response to $\boldsymbol{\mu}^{*}$, i.e. $\pi^{*}\in BR(\boldsymbol{\mu}^{*})$;
@@
Previous works have shown that, in tabular settings, it is possible for a finite population of decentralised agents (each of which is permitted to have a distinct population-independent policy $\pi^{i}$) to learn the MFNE using only the empirical distribution $\hat{\mu}_{t}$, rather than the exactly calculated infinite flow $\boldsymbol{\mu}$ Yardim et al. (2023); Benjamin and Abate (2023). This MFNE may be the goal in itself, or it can in turn serve as an approximate NE for the harder-to-solve game involving the finite population. In this work we provide algorithms to perform this process in non-tabular and population-dependent settings, and demonstrate them empirically.
@@
Instead of performing the computationally expensive process of finding a $BR$ at each iteration, we can use a form of policy iteration for MFGs called Online Mirror Descent (OMD). This involves beginning with an initial policy $\pi_{0}$, and then at each iteration $k$, evaluating the current policy $\pi_{k}$ with respect to its induced mean-field flow $\boldsymbol{\mu}=I(\pi_{k})$ to compute its Q-function $Q_{k+1}$. To stabilise the learning process, we then use a weighted sum over this and past Q-functions, and set $\pi_{k+1}$ to be the softmax over this weighted sum, i.e. $\pi_{k+1}(\cdot|s,\mu)=softmax\left(\frac{1}{\tau_{q}}\sum^{k}_{\kappa=0}{Q_{\kappa}}(s,\mu,\cdot)\right)$. $\tau_{q}$ is a temperature parameter that scales the entropy in Munchausen RL Vieillard et al. (2020); note that this is a different temperature to the one agents use when selecting which communicated parameters to adopt, denoted $\tau^{comm}_{k}$ (Sec. 3.2).
If the Q-function is approximated non-linearly using neural networks, it is difficult to compute this weighted sum. The so-called ‘Munchausen trick’ addresses this by computing a single Q-function that mimics the weighted sum using implicit regularisation based on the Kullback-Leibler (KL) divergence between $\pi_{k}$ and $\pi_{k+1}$ Vieillard et al. (2020). Using this reparametrisation gives Munchausen OMD (MOMD), detailed further in Sec. 3.1 Laurière et al. (2022b); Wu et al. (2024). MOMD does not bias the MFNE, and has the same convergence guarantees as OMD Hadikhanloo (2017); Perolat et al. (2021); Wu et al. (2024).
@@ 2.3 Networks
Definition 6 (Time-varying network).
@@
One of these networks $\mathcal{G}^{comm}_{t}$ defines which agents can communicate information to each other at time $t$. The second network $\mathcal{G}^{obs}_{t}$ is a graph defining which agents can observe each other’s states, which we use in general settings for estimating the mean-field distribution from local information. The structure of the two networks may be identical (e.g. if embodied agents can both observe the position (state) of, and exchange information with, other agents within a certain physical distance from themselves), or different (e.g. if agents can observe the positions of nearby agents, but only exchange information with agents by which they are linked via satellite, which may connect agents over long distances).
@@
Definition 7 (Time-varying state-visibility graph).
@@
0:  initial states {$s^{i}_{0}$}${}_{i=1}^{N}$; $t\leftarrow 0$
1:  $\forall i$ : Randomly initialise parameters $\theta^{i}_{0}$ of Q-networks $\check{Q}_{\theta^{i}_{0}}(o,\cdot)$, and set $\pi^{i}_{0}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{0}}(o,\cdot)\right)(a)$
@@
14:     $\forall i$ : $\pi^{i}_{k+1}(a|o)\leftarrow$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k+1}}(o,\cdot)\right)(a)$
15:     $\forall i:$ $\sigma^{i}_{k+1}\leftarrow 0$
@@
17:        Take step $\forall i:a^{i}_{t}\sim\pi^{i}_{k}(\cdot|o^{i}_{t}),r^{i}_{t}=R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t}),s^{i}_{t+1}\sim P(\cdot|s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})$
18:        $\forall i:$ $\sigma^{i}_{k+1}=\sigma^{i}_{k+1}+\gamma^{e}\cdot r^{i}_{t}$
@@
23:        $\forall i:J^{i}_{t}\leftarrow\{j\in\mathcal{N}:(i,j)\in\mathcal{E}^{comm}_{t}$}
@@
29:  return  policies $\pi^{i}_{K}$}${}_{i=1}^{N}$
3 Learning and policy improvement
3.1 Q-network and update
Lines 1-14 of our novel Alg. 1 contain the core Q-function/policy update method. Agent $i$ has a neural network parametrised by $\theta^{i}_{k}$ to approximate its Q-function: $\check{Q}_{\theta^{i}_{k}}(o,\cdot)$. The agent’s policy is given by $\pi_{\theta^{i}_{k}}(a|o)=$ softmax$\left(\frac{1}{\tau_{q}}\check{Q}_{\theta^{i}_{k}}(o,\cdot)\right)(a)$. We denote the policy $\pi^{i}_{k}(a|o)$ for simplicity when appropriate. Each agent maintains a buffer (of size $M$) of collected transitions of the form $\left(o^{i}_{t},a^{i}_{t},r^{i}_{t},o^{i}_{t+1}\right)$. At each iteration $k$, they empty their buffer (Line 3) before collecting $M$ new transitions (Lines 4-7); each decentralised agent $i$ then trains its Q-network $\check{Q}_{\theta^{i}_{k}}$ via $L$ training updates as follows (Lines 8-12). For training purposes, $i$ also maintains a target network $\check{Q}_{\theta^{i,^{\prime}}_{k,l}}$ with the same architecture but parameters $\theta^{i,^{\prime}}_{k,l}$ copied from $\theta^{i}_{k,l}$ less regularly than $\theta^{i}_{k,l}$ themselves are updated, i.e. only every $\nu$ learning iterations (Line 11). At each iteration $l$, the agent samples a random batch $B^{i}_{k,l}$ of $|B|$ transitions from its buffer (Line 9), and trains its neural network via stochastic gradient descent to minimise the empirical loss (Def. 8, Line 10). For $cl<0$, $[\cdot]^{0}_{cl}$ is a clipping function used in Munchausen RL to prevent numerical issues if the policy is too close to deterministic, as the log-policy term is otherwise unbounded Vieillard et al. (2020); Wu et al. (2024):
Definition 8 (Empirical loss for Q-network).
This is given by